A hunter aims directly at a target(on the same level) 170 m away. If the bullet leaves the gun at a speed of 40 m/s, by how much will it miss the target?
Answer by t0hierry(194)

It has to miss because of the weight of the bullet. The bullet's equation of motion is a parabola. Acceleration in y, constant velocity in x. The question is now: given the time it takes to reach 170 meters, how much did the bullet fall?
At 40 m/s, it takes 170/40 = 4.25 sec to reach as far as 170 meters. In that time, the acceleration due to gravity along the y axis gives us (Newton's law)
y = 1/2 g t^2 + v_y *t + y_0
We can take y_0 = 0 and we know that v_y_0 = 0 (no initial vertical velocity)
y = 1/2 9.81 * (4.25)^2 = 88.6 meters.
That's a big miss. Can you check that the numbers are right?